# LLMs Process Lists With General Filter Heads

### v1

@@ Appendix D Vector Algebra with Predicate Representation
We explore the geometric properties of the predicate representation $q_{\mathrm{src}}$ by examining its behavior under vector arithmetic operations. Specifically, we investigate when we compose two predicates (find the fruit and find the vehicle) by adding their corresponding $q_{\mathrm{src}}$ vectors, does this resulting vector represent a meaningful combination of the two predicates?
@@
If we add the $q_{\mathrm{src}}$ vectors of two sources prompts with different predicates, $p_{\mathrm{src1}}=\mathbb{P}(\mathcal{C}_{1},\psi_{1}=\textit{is\_fruit})$ and $p_{\mathrm{src2}}=\mathbb{P}(\mathcal{C}_{2},\psi_{2}=\textit{is\_vehicle})$, we find that the resulting vector $q_{\mathrm{composed}}=q_{\mathrm{src1}}+q_{\mathrm{src2}}$ can be used on a destination prompt $p_{\mathrm{dest}}=\mathbb{P}(\mathcal{C}_{3},\psi_{3}\notin\{\textit{is\_fruit},\textit{is\_vehicle}\})$ to execute the disjunction of the two predicates (i.e., find the fruit or vehicle) in $\mathcal{C}_{3}$. The setup is illustrated in Figure˜14 with an example from the SelectOne task.
We conduct this experiment for the SelectOne task. We compose $q_{\mathrm{composed}}$ with two prompts and curate $p_{\text{dest}}$ such that there is only item $c_{\mathrm{targ}}$ in $\mathcal{C}_{\mathrm{dest}}$ that satisfies the composed predicate $\psi_{\mathrm{src1}}\cup\psi_{\mathrm{src2}}$. We patch $q_{\mathrm{composed}}$ to the destination run and consider the intervention to be successful if the LM thinks $c_{\mathrm{targ}}$ is the most probable item. We achieve a causality score of $0.6523$ ($334$ out of $512$) with logit of $c_{\mathrm{targ}}$ increased by $6.75\pm 3.94$ after the intervention.
@@ E.1 Do these heads merely match with a pre-computed answer?
A more subtle concern is whether these heads encode abstract predicates or simply store concrete answers. For instance, when the source prompt asks to find the fruit with answer Plum, does the query state encode the predicate find the fruit or the specific item Plum? In the patched run, Plum representation would naturally show higher similarity to the representaion of Apple (another fruit) than to Watch (a non-fruit), potentially explaining the selective behavior we observe in the attention pattern after patching $q_{\mathrm{src}}$ in the patched run.
@@ Appendix F Averaging to remove the order id
When examining cases where our intervention fails, we find that the model sometimes selects the item at the same position as the original answer in the source prompt. Figure˜8 also shows elevated scores for items matching the source answer’s position $c_{\text{oid}}$ in critical layers, although not as high as $c_{\text{targ}}$. This suggests that the LM also encodes the positional information or order ID (Feng & Steinhardt, 2024; Prakash et al., 2025) of $c_{\text{src}}$ alongside the predicate. A probable explanation for this is: as these filter heads are distributed across a range of layers, patching $q_{\mathrm{src}}$ from filter heads in later layers may bring over specific contribution from the filter heads in earlier layers.
Figure 15: Averaging $q_{\text{src}}$ to remove the order ID information. This simple trick improves the causality scores by 7 to 10 points across the board. Causality scores presented for 512 examples from SelectOne task.
To isolate the predicate signal from this positional bias, we use a simple trick: averaging the query states across multiple source prompts. We produce $n$ variations of the same source prompt $p_{\text{src}}=\mathbb{P}(\mathcal{C}_{\text{src}},\psi_{\text{src}})$ by changing the index of the correct answer $c_{\text{src}}$. Figure˜15 illustrates this idea. This trick improves our causality scores, also shown in Figure˜15.
@@
Landmark in Country
0.576(295/512)0.576\,(295/512)
0.670(343/512)0.670\,(343/512)
Recall which country these
@@
Appendix H Decomposing $\text{W}_{QK}$ into Smaller Functional Units with SVD
We investigate if attention heads are too large of a model unit for analyzing the filtering task and if we can be more surgical by decomposing heads into smaller functional units. We take inspiration from Merullo et al. (2024b) and attempt locating low-rank subspaces of the $QK$ matrix necessary to execute the filtering task.
@@ Appendix H Decomposing $\text{W}_{QK}$ into Smaller Functional Units with SVD
In attention heads, the query projection $\text{W}_{Q}$ and the key projection $\text{W}_{K}$ read from the residual stream latent $h_{t}$ to determine the attention pattern, while the $OV$ projection $\text{W}_{OV}$ writes back to the residual stream. Equation˜1 shows how the attention distribution from a query token $t$ is calculated but for simplicity it hides some technical details, such as a pre-layernorm is applied on $h^{\ell-1}$ before $\text{W}_{K}$ and $\text{W}_{Q}$ are applied, and RoPE positional encoding (Su et al., 2024) is applied on the resulting query state and key states before the attention score is calculated. If we do not ignore these details the pre-softmax attention score from the $q^{th}$ token to the $k^{th}$ token is calculated as.
Figure 16: Distribution of critical SVD components of $\text{W}_{QK}$ across attention heads in Llama-70B. Middle layers show higher component counts, aligning with filter head locations.
Table 7: Causality of the SVD components (without averaging trick) on the SelectOne task. Components located using object-type selection (e.g. find the fruit) generalize to person identification (e.g. find the actor). Both in-distribution and out-of-distribution causality were calculated using 512 examples. Excluding the components in later layers significantly improves performance in both in-distribution and out-of-distribution validation sets.
Attention Heads Object-type Profession
All $0.627$ $0.420$
Heads in $\ell<52$ $0.811$ $0.773$
Only filter heads $0.766$ $0.752$
$a_{qk}=\text{RoPE}\big(\tilde{h}_{q}\text{W}_{Q}\big)\;\Big(\text{RoPE}\big(\tilde{h}_{k}\text{W}_{K}\big)\Big)^{T}$ $,\;\;\;\text{where $\tilde{h}=\text{LayerNorm\big(h\big)}$}$ (7)
When we ignore the positional encoding (i.e. setting $\text{RoPE}=I$), we can combine $\text{W}_{Q}$ and $\text{W}_{K}$ projections in to a single bilinear form.
$a_{qk}$ $=\tilde{h}_{q}\text{W}_{Q}\;\Big(\tilde{h}_{k}\text{W}_{K}\Big)^{T}$ $=\tilde{h}_{q}\,\text{W}_{QK}\,\tilde{h}_{k}^{T}$ (8)
Note that although $\text{W}_{QK}$ is a $d_{\text{embd}}\times d_{\text{embd}}$ matrix it has a rank of $d_{\text{head}}$. This is because $\text{W}_{Q}$ and $\text{W}_{K}$ both have a rank of $d_{\text{head}}$. To get the attention distribution from the $q^{th}$ token the attention score $a_{{qk}}$ is calculated for all $k\leq q$, which is then divided by a scaling factor $\sqrt{d_{\text{head}}}$ before applying softmax (see Equation˜1).
@@
Note that $u_{\text{patch}}$ is a $d_{\text{head}}$ vector where the $i^{th}$ value is the same as the $i^{th}$ value of $u_{\text{src}}$ when $\text{mask}_{i}$ is on, and the same as the $i^{th}$ value of $u_{\text{dest}}$ when $\text{mask}_{i}$ is off. Recalculating $a_{qk}$ in the patched run ensures that only the information encoded in the selected subspace is transferred from the source prompt while the information in other components is preserved. Similar to the method detailed in Section˜2 we learn this mask for all heads.
Figure˜16 shows that middle-layer attention heads contain more components relevant to the filtering task, consistent with where we find the filter heads to be located. Interestingly, while our optimization identifies components in later layers, excluding these actually improves causality (Table˜7), suggesting they may introduce noise rather than contributing information relevant for filtering. Most notably, using only the SVD components of previously identified filter heads achieves high causality (0.766 for in-distribution, 0.752 for out-of-distribution), demonstrating that predicate information is concentrated in low-rank subspaces within specialized heads. This finding indicate that filter heads implement filtering through coordinated interactions between low-rank subspaces of the query and key states, rather than using their full representational capacity.
Appendix I What do the Filter Heads write to the residual stream?
@@ 2. H2:
To distinguish between these hypotheses, we design a controlled causal mediation analysis by patching the output of the filter heads (output of OV projection). We create source and destination prompts with different predicate but the same collection of items presented in different order (see Figure˜17 left). This ensures that $c_{\text{src}}$, the correct answer of the source prompt remain a valid option in both prompts, enabling us to test H1. While shuffling we make sure that the order of $c_{\text{src}}$ is different so that we can test H2. We choose an MCQ format and ensure that the MCQ labels differs between the prompts to avoid confounding effects of this intervention.
@@
We patch the output (OV contribution to residual stream) of 79 filter heads identified with the SelectOne task (not MCQ). The results plotted in Figure˜17 (right) reveal nuanced insights:
@@
Although our findings from this patching experiment are somewhat inconclusive, they suggest that filter heads write both what items are filtered and where they are located. This dual encoding strategy aligns with our earlier finding of multiple implementation pathways. This redundancy may enable the LM to flexibly use either information or both of them in parallel.
Appendix J Distinguishing Filter Heads from Other Specialized Attn Heads
We compare filter heads with two previously characterized attention head types of establish their distinctiveness. Function Vector (FV) heads, identified in Todd et al. (2024), encode a compact representation of the task demonstrated in few-shot examples. And, Concept Induction (CI) heads, identified in Feucht et al. (2025), encode the concept or lexical unit in multi-token words.
Following the identification procedures detailed in the respective papers, we locate FV and CI heads in Llama-70B, and compare their distribution with filter heads in Figure˜18. Interestingly, although both FV heads and filter heads are concentrated in the middle layers, only 7 (out of 79) heads overlap. This separation suggests that these head types play parallel, complementary roles rather than performing redundant operations.
Figure 18: Distribution of specialized attention heads in Llama-70B. Both Filter and FV heads are concentrated in the middle layers, but there are only 7 overlapping heads. CI heads exhibit weaker concentration in the middle layers, with minimal overlap with filter heads (only one head). For a fairer comparison with choose 79 heads from each head type.
@@
Appendix K Replicating experiments in Gemma-27B
@@
K.1 Location of the Filter Heads
Figure 19: Location of filter heads in Gemma-27B-it. Similar to Figure˜13, (a) shows the individual FilterScore for each head: how much they attend to the correct option over others. (b) shows the indirect effect: how much patching $q_{\mathrm{src}}$ from a single head promote the predicate target. The filter heads identified with the method detailed in Section˜2.3 are marked with black borders.
K.2 Within Task Generalizability
@@
K.3 Across Task Generalizability
@@
K.4 Dual Implementation in Question Before vs Question After
@@
Appendix L Limitations
@@
We examined only six filter-reduce tasks, which may not capture the full diversity of filtering strategies employed by LLMs. Even within our six tasks we identified that the filter heads do not show high causality in the CheckPresence task, indicating that the LM uses alternate mechanism for certain filtering operations. The consistent prompt templates we used enable us to scale up our controlled experiment setup, but they may have biased us towards specific computational strategies inside the LM. LMs may adapt their filtering approach based on what information is available in ways that our limited task set and prompting strategies cannot fully reveal.
@@
We identified filter heads in Llama-70B and Gemma-27B models. The fact that we were able to identify similar mechanisms in two models of different sizes, from different families, trained on different datasets echos the idea of Evolutionary Convergence from Morris (2006): distantly related organisms (e.g. vertibrate and cephalopods) independently evolve similar adaptations (e.g. eyes) in response to similar environmental pressure.
However, such convergence is not guaranteed across all LMs. Notably, findings from Zhong et al. (2023) suggest that identical architectures trained on different datasets can potentially develop different implementations for the same operation. LMs from other families may develop a mechanism that does not make use of such filter heads. Additionally, we restrict our analysis to fairly larger LMs. It is possible that smaller LMs, where parameter constraints might enforce higher head-level superposition, entangle the predicate representation may get entangled with other unrelated stuff in the query states. Therefore, we might not see the distinct head-level causal role of filter heads we get in larger LMs.
@@
Our tasks are designed such that we can determine whether an answer is correct based on what the LM predicts as the next token. While curating a prompt we ensure that none of the items share a first token with each other. In all of our experiments we also ensure that the answer of the source prompt, destination prompt, and the target answer of the transported predicate are all different, in order to ensure that patching does not merely copy over the answer. However, this choice of validating with only the first predicted token has potentially restricted us from exploring other tasks that have a similar filter-reduce pattern of operation. Most of the causality scores we report in this paper were calculated on a single trial with a 512 examples sampled randomly. It is possible that these scores will change slightly on a different trial.
Appendix M Qualitative Attention Patterns of Filter Heads
M.1 Typical Filter Head Behavior
The following are cases where the model clearly attends to the option which corresponds to the given predicate.
@@
M.2 Filter Heads are Sensitive to Semantic Information
Filter heads show poor causality when filtering requires reasoning with non-semantic information such as phonological similarity and letter counting. The aggregated attention pattern also do not show the distinctive filtering behavior.
@@
We curate a paragraph/free-form text that mix factual and false statements about a topic. We break the text into sentences and then append a question asking the LM to identify the false information. We visualize the aggregated attention pattern of the filter heads, which focus on the last token of the false statements. To aid visualization we draw a red line underlining the sentences whose last token gets attention score exceeding a preset threshold.
We apply the same approach to sentiment analysis using movie reviews containing both positive and negative sentences.
@@
Appendix N Filter Heads in Llama-70B

### v3

@@ Appendix D Vector Algebra with Predicate Representation
We explore the geometric properties of the predicate representation $q_{\mathrm{src}}$ by examining its behavior under vector arithmetic operations. Specifically, we investigate whether, when we compose two predicates (find the fruit and find the vehicle) by adding their corresponding $q_{\mathrm{src}}$ vectors, this resulting vector represents a meaningful combination of the two predicates.
@@
If we add the $q_{\mathrm{src}}$ vectors of two source prompts with different predicates, $p_{\mathrm{src1}}=\mathbb{P}(\mathcal{C}_{1},\psi_{1}=\textit{is\_fruit})$ and $p_{\mathrm{src2}}=\mathbb{P}(\mathcal{C}_{2},\psi_{2}=\textit{is\_vehicle})$, we find that the resulting vector $q_{\mathrm{composed}}=q_{\mathrm{src1}}+q_{\mathrm{src2}}$ can be used on a destination prompt $p_{\mathrm{dest}}=\mathbb{P}(\mathcal{C}_{3},\psi_{3}\notin\{\textit{is\_fruit},\textit{is\_vehicle}\})$ to execute the disjunction of the two predicates (i.e., find the fruit or vehicle) in $\mathcal{C}_{3}$. The setup is illustrated in Figure˜14 with an example from the SelectOne task.
We conduct this experiment for the SelectOne task. We compose $q_{\mathrm{composed}}$ with two prompts and curate $p_{\text{dest}}$ such that there is only one item $c_{\mathrm{targ}}$ in $\mathcal{C}_{\mathrm{dest}}$ that satisfies the composed predicate $\psi_{\mathrm{src1}}\cup\psi_{\mathrm{src2}}$. We patch $q_{\mathrm{composed}}$ to the destination run and consider the intervention to be successful if the LM thinks $c_{\mathrm{targ}}$ is the most probable item. We achieve a causality score of $0.6523$ ($334$ out of $512$) with logit of $c_{\mathrm{targ}}$ increased by $6.75\pm 3.94$ after the intervention.
@@ E.1 Do these heads merely match with a pre-computed answer?
A more subtle concern is whether these heads encode abstract predicates or simply store concrete answers. For instance, when the source prompt asks to find the fruit with the answer being Plum, does the query state encode the predicate find the fruit or the specific item Plum? In the patched run, the Plum representation would naturally show higher similarity to the representation of Apple (another fruit) than to Watch (a non-fruit), potentially explaining the selective behavior we observe in the attention pattern after patching $q_{\mathrm{src}}$ in the patched run.
@@ Appendix F Averaging to remove the order id
When we examine the cases where our intervention fails, we find that the model sometimes selects the item at the same position as the original answer in the source prompt. Figure˜8 also shows elevated scores for items that match the position of the source answer $c_{\text{oid}}$ in critical layers, although not as high as $c_{\text{targ}}$. This suggests that the LM also encodes the positional information or order ID (Feng & Steinhardt, 2024; Prakash et al., 2025) of $c_{\text{src}}$ alongside the predicate. A probable explanation for this is that, since these filter heads are distributed across a range of layers, patching $q_{\mathrm{src}}$ from filter heads in later layers may bring over specific contributions from filter heads in earlier layers.
Figure 15: Averaging $q_{\text{src}}$ to remove the order ID information. This simple trick improves the causality scores by 7 to 10 points across the board. Causality scores presented for 512 examples from the SelectOne task.
To isolate the predicate signal from this positional bias, we use a simple trick: averaging the query states across multiple source prompts. We produce $n$ variations of the same source prompt $p_{\text{src}}=\mathbb{P}(\mathcal{C}_{\text{src}},\psi_{\text{src}})$ by changing the index of the correct answer $c_{\text{src}}$. Figure˜15 illustrates this idea. This simple averaging improves our causality scores.
@@
Appendix H Ablation on Design Choices
In this section we conduct ablation studies on two design choices we made in this work.
Sparsity Regularizer.
To identify the minimal set of heads necessary for filtering, we incorporate the L1 norm of the $\mathrm{mask}$ in our optimization objective. This sparsity constraint ensures we select only the most influential heads while filtering out noisy ones.
[…]
The target loss maximizes the logit of $c_{\mathrm{targ}}$ (the item satisfying the source predicate $\psi_{\mathrm{src}}$), while suppressing the other items in $\mathcal{C}_{\mathrm{dest}}$.
Figure 16: Sparsity-performance tradeoff. Optimal causality achieved with 79 heads identified with $\lambda=0.02$. Lower values of $\lambda$ includes noisy heads while larger values over-prune critical heads.
The sparsity penalty, modulated by the hyperparameter $\lambda$, pushes the mask towards zero so that fewer heads are selected.
We find that selecting the right sparsity coefficient is crucial to achieve higher causality. Through a systematic search across $\lambda\in(0,1]$, while controlling for other variables (training data, other hyperparameters) of the optimization process, we find a clear sweet spot at $\lambda=0.02$. In Figure˜16 we report the causality of identified heads across format, where the source prompt is from SelectOne and the destination prompt is from SelectOne-MCQ. While evaluating, we do not employ the averaging trick (Appendix˜F), and transfer head states at only the last token.
This tradeoff is intuitive: insufficient regularization allows noisy irrelevant heads into the mask, reducing generalization performance. And excessive regularization ends up pruning essential heads, degrading causality. We have identified this optimal value of $\lambda$ with Llama-70B but use the same value for all the models we investigate in this paper.
Last two tokens for evaluation.
[…]
Table 10: Evaluation by patching only the last token for SelectOne task.
Semantic Type Causality $\Delta\mathrm{logit}$
Object 0.819(839/1024)0.819\;(839/1024) $8.12\pm 2.93$
Person Profession 0.842(431/512)0.842\;(431/512) $7.31\pm 3.04$
Person Nationality (+ Priming Prefix) 0.629(322/512)0.629\;(322/512) $6.28\pm 3.75$
Landmark in country (+ Priming Prefix) 0.678(347/512)0.678\;(347/512) $7.05\pm 3.20$
For the SelectOne-Object task, single-token patching achieves a causality of $0.819$ — only slightly lower than the $0.863$ we report in the main text with two-token patching (Figure˜2). This suggests that while preventing information leakage helps, it is not as critical as we initially anticipated. The filter heads remain robustly causal even with this simpler intervention.
Appendix I Decomposing $\text{W}_{\mathrm{QK}}$ into Smaller Functional Units with SVD
We investigate whether attention heads are too large of a model unit for analyzing the filtering task and if we can be more surgical by decomposing heads into smaller functional units. We take inspiration from Merullo et al. (2024b) and attempt to locate low-rank subspaces of the $QK$ matrix necessary to perform the filtering task.
@@ Appendix I Decomposing $\text{W}_{\mathrm{QK}}$ into Smaller Functional Units with SVD
In attention heads, the query projection $\text{W}_{Q}$ and the key projection $\text{W}_{K}$ read from the residual stream latent $h_{t}$ to determine the attention pattern, while the $OV$ projection $\text{W}_{OV}$ writes back to the residual stream. Equation˜1 shows how the attention distribution from a query token $t$ is calculated, but for simplicity it hides some technical details, such as the fact that a pre-layernorm is applied on $h^{\ell-1}$ before $\text{W}_{K}$ and $\text{W}_{Q}$ are applied, and RoPE positional encoding (Su et al., 2024) is applied on the resulting query state and key states before the attention score is calculated. If we do not ignore these details the pre-softmax attention score from the $q^{th}$ token to the $k^{th}$ token is calculated as
$a_{qk}=\mathrm{RoPE}\big(\tilde{h}_{q}\text{W}_{Q}\big)\;\Big(\mathrm{RoPE}\big(\tilde{h}_{k}\text{W}_{K}\big)\Big)^{T}$ $,\;\;\;\text{where $\tilde{h}=\mathrm{LayerNorm\big(h\big)}$}$ (8)
When we ignore the positional encoding (i.e. setting $\mathrm{RoPE}=I$), we can combine $\text{W}_{Q}$ and $\text{W}_{K}$ projections in to a single bilinear form.
$a_{qk}$ $=\tilde{h}_{q}\mathrm{W}_{Q}\;\Big(\tilde{h}_{k}\mathrm{W}_{K}\Big)^{T}=\tilde{h}_{q}\,\mathrm{W}_{Q}\mathrm{W}_{K}^{T}\,\tilde{h}_{k}^{T}=\tilde{h}_{q}\,\mathrm{W}_{QK}\,\tilde{h}_{k}^{T}$ (9)
Note that although $\mathrm{W}_{QK}$ is a $d_{\mathrm{embd}}\times d_{\mathrm{embd}}$ matrix it has a rank of $d_{\text{head}}$. This is because $\text{W}_{Q}$ and $\text{W}_{K}$ both have a rank of $d_{\text{head}}$. To get the attention distribution from the $q^{th}$ token the attention score $a_{{qk}}$ is calculated for all $k\leq q$, which is then divided by a scaling factor $\sqrt{d_{\text{head}}}$ before applying softmax (see Equation˜1).
@@
Figure 17: Distribution of critical SVD components of $\text{W}_{QK}$ across attention heads in Llama-70B. Middle layers show higher component counts, which aligns with filter head locations.
Table 11: Causality of the SVD components (without the averaging trick) on the SelectOne task. Components located using object-type selection (e.g. find the fruit) generalize to person identification (e.g. find the actor). Both in-distribution and out-of-distribution causality were calculated using 512 examples. Excluding the components in later layers significantly improves performance in both the in-distribution and out-of-distribution validation sets.
Attention Heads Object-type Profession
All $0.627$ $0.420$
Heads in $\ell<52$ $0.811$ $0.773$
Only filter heads $0.766$ $0.752$
Note that $u_{\text{patch}}$ is a $d_{\text{head}}$ vector where the $i^{th}$ value is the same as the $i^{th}$ value of $u_{\text{src}}$ when $\text{mask}_{i}$ is on, and the same as the $i^{th}$ value of $u_{\text{dest}}$ when $\text{mask}_{i}$ is off. Recalculating $a_{qk}$ in the patched run ensures that only the information encoded in the selected subspace is transferred from the source prompt while the information in other components is preserved. Similarly to the method detailed in Section˜2, we learn this mask for all heads.
Figure˜17 shows that middle-layer attention heads contain more components relevant to the filtering task, which is consistent with where we locate filter heads. Interestingly, while our optimization identifies components in later layers, excluding these actually improves causality (Table˜11), which suggests that they may introduce noise rather than information relevant for filtering. Most notably, using only the SVD components of previously identified filter heads achieves high causality (0.766 for in-distribution, 0.752 for out-of-distribution), demonstrating that predicate information is concentrated in low-rank subspaces within specialized heads. This finding indicates that filter heads implement filtering through coordinated interactions between low-rank subspaces of the query and key states, rather than using their full representational capacity.
Appendix J What do the Filter Heads write to the residual stream?
@@ 2. H2:
To distinguish between these hypotheses, we design a controlled causal mediation analysis by patching the output of the filter heads (output of the OV projection). We create source and destination prompts with different predicates but the same collection of items presented in different orders (see Figure˜18 left). This ensures that $c_{\text{src}}$, the correct answer of the source prompt remains a valid option in both prompts, enabling us to test H1. While shuffling, we make sure that the order of $c_{\text{src}}$ is different so that we can test H2. We choose an MCQ format and ensure that the MCQ labels differ between the prompts to avoid confounding effects of this intervention.
@@
We patch the output (OV contribution to the residual stream) of 79 filter heads identified with the SelectOne task (not MCQ). The results plotted in Figure˜18 (right) reveal nuanced insights:
@@
Although our findings from this patching experiment are somewhat inconclusive, they suggest that filter heads write both what items are filtered and where they are located. This dual encoding strategy aligns with our earlier finding of multiple implementation pathways. This redundancy may enable the LM to flexibly use either piece of information, or both of them in parallel.
Appendix K Distinguishing Filter Heads from Other Specialized Attn Heads
We compare filter heads with two previously characterized attention head types to establish their distinctiveness. Function Vector (FV) heads, identified in Todd et al. (2024), encode a compact representation of the task demonstrated in few-shot examples. And, Concept Induction (CI) heads, identified in Feucht et al. (2025), encode the concept or lexical unit in multi-token words.
Following the identification procedures detailed in the respective papers, we locate FV and CI heads in Llama-70B. We use publicly available dataset and source code from Todd et al. (2024) and Feucht et al. (2025). In Figure˜19 we plot the distribution of these heads with filter heads identified with SelectOne task. Interestingly, although both FV heads and filter heads are concentrated in the middle layers, only 7 (out of 79) heads overlap. This separation suggests that these head types play parallel, complementary roles in LLM computation.
Figure 19: Distribution of specialized attention heads in Llama-70B. Both Filter and FV heads are concentrated in the middle layers, but there are only 7 overlapping heads. CI heads exhibit weaker concentration in the middle layers, with minimal overlap with filter heads (only one head). For a fairer comparison, we choose 79 heads from each head type.
@@
Appendix L Replicating experiments in Gemma-27B
@@
L.1 Location of the Filter Heads
Figure 20: Location of filter heads in Gemma-27B-it. Similar to Figure˜13, (a) shows the individual FilterScore for each head: how much they attend to the correct option over others. (b) shows the indirect effect: how much patching $q_{\mathrm{src}}$ from a single head promotes the predicate target. The filter heads identified with the method detailed in Section˜2.3 are marked with black borders.
L.2 Within Task Generalizability
@@
L.3 Across Task Generalizability
@@
L.4 Dual Implementation in Question Before vs Question After
@@
Appendix M Filter Heads in Smaller LMs: Llama-8B and Gemma-9B
We investigate whether filter heads also emerge in smaller models from the same families: Llama-8B and Gemma-9B. We find that filter heads do exist in these smaller models, though they show slightly weaker cross-task generalization compared to their larger counterparts (Llama-70B and Gemma-27B). We hypothesize that this degradation likely because of the increased parameter constraints in smaller models individual attention heads are more inclined to serve multiple computational roles, causing the predicate signal to become more entangled with other information. However, these results hint that the filter head mechanism is architecturally general in Llama and Gemma family of LMs, but its abstraction and portability improve with model scale.
Table 15: Causality of filter heads on SelectOne tasks. Using object-type filtering queries we locate 42 heads in Gemma-9B and 47 heads in Llama-8B. We check if the same heads generalize to semantically different predicates like profession identification. Compare with Llama-70B in Figure˜2 and Gemma-27B in Section˜L.2.
Semantic Type Model
Gemma-9B (42 Heads) Llama-8B (39 Heads)
Object 0.856(438/512)0.856\;(438/512) 0.826(423/512)0.826\;(423/512)
Person Profession 0.828(424/512)0.828\;(424/512) 0.756(387/512)0.756\;(387/512)
Person Nationality (+ Priming Prefix) 0.387(198/512)0.387\;(198/512) $0.525(269/512)$
Landmark in country (+ Priming Prefix) 0.492(252/512)0.492\;(252/512) 0.492(252/512)0.492\;(252/512)
Figure 24: Generalization of filter heads across different tasks on Llama-8B and Gemma-9B. Compare with Figure˜3 for Llama-70B and Figure˜21 for Gemma-27B.
Appendix N Usage of LLMs
Proprietery LLMs services with black-box access such as claude.ai and gemini.google.com were used as a general purpose assist tool, which is allowed as per the ICLR 2026 author guideline. We have used such LLMs to polish some of the writings in this paper. We have also used LLMs to get more items for our dataset and translating the prompts to other languages.
Appendix O Limitations
@@
We examined only six filter-reduce tasks, which may not capture the full diversity of filtering strategies employed by LLMs. Even within our six tasks we identified that the filter heads do not show high causality in the CheckPresence task, indicating that the LM uses alternate mechanisms for certain filtering operations. The consistent prompt templates we used enabled us to scale up our controlled experiment setup, but they may have biased us towards specific computational strategies inside the LM. LMs may adapt their filtering approach based on what information is available in ways that our limited task set and prompting strategies cannot fully reveal.
@@
We identified filter heads in Llama-70B, Llama-8B, Gemma-27B, and Gemma-9B models. The fact that we were able to identify similar mechanisms in four models of different sizes, from different families, trained on different datasets echoes the idea of Evolutionary Convergence from Morris (2006): distantly related organisms (e.g. vertebrate and cephalopods) independently evolve similar adaptations (e.g. eyes) in response to similar environmental pressure.
However, such convergence is not guaranteed across all LMs. Notably, findings from Zhong et al. (2023) suggest that identical architectures trained on different datasets can potentially develop different implementations for the same operation. LMs from other families may develop a mechanism that does not make use of such filter heads. Additionally, we restrict our analysis to fairly larger LMs. It is possible that smaller LMs, where parameter constraints might enforce higher head-level superposition, the predicate representation may get entangled with other unrelated stuff in the query states. Therefore, we might not see the distinct head-level causal role of filter heads we get in larger LMs.
@@
Our tasks are designed such that we can determine whether an answer is correct based on what the LM predicts as the next token. While curating a prompt we ensure that none of the items share a first token with each other. In all of our experiments we also ensure that the answer of the source prompt, destination prompt, and the target answer of the transported predicate are all different, in order to ensure that patching does not merely copy over the answer. However, this choice of validating with only the first predicted token has potentially restricted us from exploring other tasks that have a similar filter-reduce pattern of operation. Most of the causality scores we report in this paper were calculated on a single trial with 512 examples sampled randomly. It is possible that these scores would change slightly on a different trial.
Appendix P Qualitative Attention Patterns of Filter Heads
In this section we provide the aggregated attention pattern of the filter heads in different examples. We use the same 79 filter heads that we identified in Llama-70B for the SelectOne task. The query token is the last token in all cases.
P.1 Typical Filter Head Behavior on Tasks We Investigate
Filter heads exhibit a distinct attention pattern: when given a filtering task, they selectively focus their attention on items that satisfy the filtering criteria. Below we visualize this behavior across the six tasks we study in the main text.
@@
P.2 Generalization Beyond Our Core Tasks
To check the generalization of filter heads beyond our six core tasks, we examine their attention patterns on additional filtering scenarios that we did not systematically evaluate in the main text. These examples include negated predicates, logical deduction, mathematical reasoning, and identifying odd-one-out items.
Across these diverse scenarios, we observe that the same 79 heads identified for the SelectOne task maintain their characteristic filtering behavior: their aggregated attention consistently focuses on items satisfying the given predicate. This supports our claim that a range of different filtering operations share a common computational circuit implemented by a consistent set of filter heads.
Not Predicate [Uncaptioned image] Logical Deduction [Uncaptioned image]
Math Reasoning [Uncaptioned image] Select Odd-One-Out [Uncaptioned image]
P.3 Failure Cases of Filter Heads
We also identify boundary cases where filter heads do not exhibit their characteristic attention patterns. These tasks require reasoning with non-semantic information—such as phonological similarity, letter counting, alphabetical ordering. This suggests that filtering operations requiring non-semantic reasoning may rely on different computational mechanisms. These alternate mechanisms may not involve the filter heads we identified or these heads may play some other role which we have not investigated within the scope of this work.
@@
We curate paragraphs/free-form texts that mix factual and false statements about a topic. We break the text into sentences and then append a question asking the LM to identify the false information. When we visualize the aggregated attention pattern of the filter heads, identified with the SelectOne task, we observe that the heads focus their attention on the last token of the sentences containing false information. We annotate the last tokens of the sentences with a black border to aid visualization.
We can apply the same approach for sentiment analysis in movie/item reviews containing both positive and negative sentences.
@@
Negative Sentiment Detector 2 [Uncaptioned image]
Appendix Q Filter Heads in Llama-70B